OpenStudy (nilankshi):

(2x-1)(2x_3)+(3x-1)(3x-2)

OpenStudy (anonymous):

(2x-1)(2x_3) =(2x)(2x)-3(2x)-1(2x)+(-1)(-3) =4x^2 -6x -2x +3 =4x^2 -8x +3 (3x-1)(3x-2) =(3x)(3x)-2(3x)-3x+(-1)(-2) =9x^2 -6x -3x +2 =9x^2 -9x +2 (2x-1)(2x_3)+(3x-1)(3x-2) =4x^2 -8x +3 + 9x^2 -9x +2 =13x^2 -17x +5

OpenStudy (nilankshi):

thzz..